OpenStudy (anonymous):

Simplify completely x^2 + 12x +35 ________________ 3x + 15

OpenStudy (anonymous):

Numerator: \[x^2+12x+35=(x+7)(x+5)\] Denominator: \[3x+15=3(x+5)\] so you have \[(x+7)(x+5)/3(x+5)\] You can cancel out the (x+5) terms on the bottom and top and you end up with: \[(x+7)/3\]

OpenStudy (anonymous):

or x/3+7/3

OpenStudy (anonymous):

thank you!!!